OpenStudy (anonymous):

The solution to x2 − 12x = 0 is {1, 12} true or false?

OpenStudy (mikurout):

false

OpenStudy (johnweldon1993):

So we have an ordered pair (1 , 12) as we know that is (x , y) so if we plug in 1 for 'x' in your equation do we get the result that y= 12? lets check \[\large x^2 - 12x = 0\] \[\large (1)^2 - 12(1) = 0\] \[\large 1 - 12 = 0\] \[\large -11 \cancel{=} 0\] so it is false

OpenStudy (mikurout):

@johnweldon1993 .(0,12) is this a solution to the above question?

OpenStudy (johnweldon1993):

No, there are only 2 solutions to this equation and they are \[\large x^2 - 12x = 0\] \[\large x^2 - 12x + 36 = 36\] \[\large (x - 6)^2 = 36\] \[\large x - 6 = \pm6\] \(\large x = 0\) or \(\large x = 12\)

OpenStudy (mikurout):

Same thing. (0,12) is the solution. x1 = 0 and x2 = 12

OpenStudy (johnweldon1993):

Or in ordered pair form (0 , 0) or (12 , 0) as you can see 'y' will always be 0 because that is the solution to a polynomial (where the parabola crosses the x-axis)

OpenStudy (mikurout):

yes , we have no y terms in our equation so y will be always zero. Thanks.